How do I write the following in simplest form?

5 3/5 to 2 1/10

Answer by Alan3354(69443)

How do I write the following in simplest form?
5 3/5 to 2 1/10
--------------
Do you mean as a ratio?
5.6/2.1 = 56/21 = 8/3, or 8 to 3